TIP: Click on subject to list as thread! ANSI
echo: sb-nasa_news
to: All
from: Hugh S. Gregory
date: 2003-02-06 18:08:00
subject: 1\19 ISS On Orbit Status 19-01-2003

This Echo is READ ONLY !   NO Un-Authorized Messages Please!
 ~~~~~~~~~~~~~~~~~~~~~~~~   ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~

ISS On-Orbit Status 1/19/03

All ISS systems continue to function nominally, except as noted 
previously or below. Ahead: Week 8 for Expedition 6.

Sunday, and 2nd weekend rest day for the crew, but...never a dull 
moment.

FE1-Nikolai Budarin, assisted by his crewmates, had a very busy 
schedule, dominated by Vozdukh troubleshooting. [After preparing for 
the checkout of the CO2 scrubber's suspected VN vacuum pump and 
demating its BITS2-12 internal telemetry connection, he was to remove 
the pump for a series of measurements of electrical resistance between 
pump connector terminals/pins with the portable ohmmeter. Depending on 
the measurement results, after consultation with MCC-M/TsUP, he then 
was to install either a new spare pump or reinstall the old pump for 
further troubleshooting ahead. If successful, Vozdukh was to be 
activated, in automatic cycle mode. The VN pump moves air from the CO2 
absorption cartridges into the station atmosphere to minimize air loss 
before vacuum regeneration of each of the two filters. The VN can 
reduce the cartridge's pressure down to 50 mmHg, and its operation is 
monitored by observing the residual pressure in the cartridge.]

Budarin also performed the weekly readings of the SVO water supply 
status counter as well as SP urine collection and pretreat assembly 
inspection and toilet flush counter for calldown to MCC-M/TsUP via 
S-band, followed by the weekly daily routine inspection of the water 
processor's BRPK air/condensate separator system.

Later, Budarin completed the Russian BIO-5 Rasteniya-2/Lada-2 
("Plants-2") plant growth payload, then terminated the 24-hr. 
regeneration cycle for adsorbent bed #2 of the BMP harmful impurities 
filtration unit. Both filter channels are now back in Purify mode.

Nikolai also performed the daily routine maintenance of the SOSh life
support system, while FE-2/SO Don Pettit prepared the daily IMS 
inventory database file for automatic update import/export.

Pettit then collected his two daily CO2 readings in Lab and SM using 
the CDMK (carbon dioxide monitor kit).

FE-1 Budarin was scheduled for a televised PFC (private family 
conference) to TsUP via K-band and S-band. PFCs were also scheduled 
for CDR Ken Bowersox and Don Pettit, via audio/S-band.

All crewmembers performed their daily physical exercise.

The crew was congratulated on an excellent R&R job yesterday, when 
they removed and replaced an O2 (oxygen) regulator in the U.S. 
Airlock. [The ground requested further clarification of an observation 
by Pettit, who, after the R&R, had mentioned a "knick" on the surface 
of the discharge line from the regulator, apparently referring to the 
secondary sealing surface of a Gammah fitting in the line.]

Questions were also uplinked to Don regarding problems he encountered 
on 1/17 with the TOCA (total organic carbon analyzer). [The questions 
are intended to help clarify the circumstances surrounding an error 
message that popped up after Pettit had set up the TOCA reagent mix, 
and which he was unable to clear up before stowing the equipment. TOCA 
is a water quality monitoring instrument for determining 
concentrations of total organic carbon, total inorganic carbon, total 
carbon, pH value and conductivity in water.]

CDR Bowersox was thanked for his assistance and inputs during I6 
processing runs of the ZCG (Zeolite crystal growth) experiment, which 
he completed yesterday with the deactivation of the ZCG furnace 
heaters. [His observations from the clear autoclaves, and suggestions 
for coalescing bubbles, were most welcome to payload scientists, who 
are now keen on seeing if fewer nucleation points indeed yields some 
of their best science yet.]

POC (Payload Operations Center) reported to Bowersox that his EXPRESS 
Rack 5 (ER5) software load on 1/17 for the ELC (ER5 laptop computer) 
and RIC (rack interface controller) was successful, but that POC 
afterwards did not see ER5 health and status properly, necessitating a 
software change. [Analysis of a command log file sent down by Sox 
showed the cause to be a software glitch (improper setting of an IP 
address) which will be corrected prior to the planned software loading 
of ER4 on 1/24. Loading of ER1 tomorrow, 1/20, will not be affected.]

MCC-M/TsUP is planning Progress M1-258 (9P) propellant transfer (fuel 
and oxidizer) to the ISS Russian segment next week (probably 1/23).

Today's targets for the CEO (crew earth observations) program were 
Manaus region, Brazil (Dynamic Event. Clear weather in the Amazon 
basin. A mapping swath of the multiple waterways of the Negro River 
[second largest on the planet] requested. The Negro dug a canyon in 
inner Amazonia during the last ice age. This canyon is now filled with 
young sediment with so many islands that the region is known locally 
as the "archipelago". Looking left for highly sensitive sunglint 
mapping views of the Negro and its poorly mapped tributaries), Buenos 
Aires, Argentina (nadir view; ESC [electronic still camera]),and 
Patagonian Glaciers (detailed views of the highly dynamic lower ends 
of glaciers coming off the Andes Mountains are requested. Looking left 
and right of track).  CEO images can be viewed at the website
http://eol.jsc.nasa.gov

ISS Orbit (as of this morning, 6:31am EST [= epoch]):  Mean altitude 
-- 389.4 km Apogee -- 393.3 km Perigee -- 385.5 km Period -- 92.34 
min.  Inclination (to Equator) -- 51.64 deg Eccentricity -- 0.0005711
Orbits per 24-hr. day -- 15.59 Altitude loss -- 170 m (mean) in last 
24 hours Revolutions since FGB/Zarya launch (Nov. '98) -- 23780

For more on ISS orbit and worldwide naked-eye visibility dates/times, 
see http://www.hq.nasa.gov/osf/station/viewing/issvis.html

 - End of File -
================

---
* Origin: SpaceBase[tm] Vancouver Canada [3 Lines] 604-473-9357 (1:153/719)
SEEN-BY: 633/267 270
@PATH: 153/719 715 7715 140/1 106/2000 1 379/1 633/267

SOURCE: echomail via fidonet.ozzmosis.com

Email questions or comments to sysop@ipingthereforeiam.com
All parts of this website painstakingly hand-crafted in the U.S.A.!
IPTIA BBS/MUD/Terminal/Game Server List, © 2025 IPTIA Consulting™.